# Targeted Immunisation Thresholds for the Contact Process on Power-Law Trees

### v1

@@ 5.2 Small truncation threshold
First we consider the $k=1$ case: if $v\sim o$, then after removing $v$ we argue via Lemma 5.1 that the expected total infectious period of $o$ is bounded by $88\lambda^{-72\rho}$. This uses the fact that the infection on $\mathscr{T}$ dominates the infection on this subtree with root degree $\mathbbm{t}-1$. During this time period $v$ simply receives infections according to a Poisson process of rate $\lambda$ and thus $m^{[v]}_{\infty}\leq 88\lambda^{1-72\rho}$ in that case.
If instead $k\geq 2$ we let $w\sim o$ be the neighbour of the root separating $o$ and $v$. We can control the quantity of interest by considering the CP-BRW with sinks $\mathcal{V}=\{o,w,v\}$. The previous argument for $k=1$ found that the mean transmission from $o$ to $w$ is
$m^{\mathcal{V},[v]}(o,w)=m^{[w]}_{\infty}\leq 88\lambda^{1-72\rho}.$
To understand the infection from $w$, let $\mathcal{Z}$ be the minimal subtree of $\mathscr{T}$ containing the three sinks $o$, $w$ and $v$. On this tree from initial infection $\mathbbm{1}_{w}$, we construct two supermartingales. The first
$M^{(o)}_{t}:=\sum_{z\in\mathcal{Z}}(2\lambda)^{\operatorname{d}(o,z)}\xi^{[o]}_{t}(z)$
is a supermartingale by the branching random walk calculation of [mountford2013metastable, Lemma 5.1]. Hence the expected number of infections of $o$ by the infection at $w$ is given by
$m^{\mathcal{V},[v]}(w,o)\leq\operatorname{\mathbb{E}}[\xi^{[o]}_{\infty}(o)]\leq\mathbb{E}\left(M^{(o)}_{\infty}\big{|}M^{(o)}_{0}\right)\leq M^{(o)}_{0}=2\lambda.$
Similarly by considering
$M^{(v)}_{t}:=\sum_{z\in\mathcal{Z}}(2\lambda)^{\operatorname{d}(v,z)}\xi^{[v]}_{t}(z)$
we can argue that $m^{\mathcal{V},[v]}(w,v)\leq(2\lambda)^{k-1}$. We can now use these three means in (3) to obtain the upper bound
$m^{[v]}_{\infty}\leq 88\lambda^{1-72\rho}\cdot(2\lambda)^{k-1}\sum_{i=0}^{\infty}\left(2\lambda\cdot 88\lambda^{1-72\rho}\right)^{i}$
For small $\rho$ and $\lambda$ the reason of the geometric series is smaller than $\lambda$, thus the sum is bounded by $1/(1-\lambda)$ and we obtain the required result by reducing again the value of $\lambda$ if necessary. ∎
The former results holds on the tree $\mathscr{T}$ which contains vertices of degree $\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor$ and only one high-degree vertex, of degree $\mathbbm{t}$. We now generalize this result to the case of a tree $\mathcal{B}$ that contains several vertices of degree $\mathbbm{t}$, but where only a small proportion $q$ of the neighbours of any given vertex has degree $\mathbbm{t}$.
@@
$m^{[v]}_{\infty}\leq\left(91\lambda\right)^{\operatorname{d}(o,v)-72\rho\,n_{\rm big}(o,v)}.$
@@
$m^{\mathcal{V},[v]}(w,u)\leq 45\lambda^{-72\rho}(2\lambda)^{d(w,u)}.$
@@
$45\lambda^{-72\rho}\left((2\lambda)^{k}\cdot q\mathbbm{t}\left(\frac{1}{8\lambda^{2}}\right)^{k-1}+\sum_{m=1}^{\infty}(2\lambda)^{2m+k}\cdot q\left(\frac{1}{8\lambda^{2}}\right)^{m+k}\right)$
$=45\lambda^{-72\rho}(4\lambda)^{-k}\left(q\mathbbm{t}\cdot 8\lambda^{2}+q\right)\leq 45\lambda^{-72\rho}(4\lambda)^{-k}\left(9\lambda^{2}q\mathbbm{t}\right).$
@@
$\sum_{m=1}^{\infty}q\left(\frac{1}{8\lambda^{2}}\right)^{m}\cdot 45\lambda^{-72\rho}(2\lambda)^{2m}=45q\lambda^{-72\rho}.$
@@
Recall now the notation $\left(\phi^{\mathcal{V},[v]}_{n}\right)_{n\in\mathbb{N}}$ from Definition 2.4 for the BRW on the set of sinks with absorption at $v$, and define
@@
for some parameter $\alpha$. Putting the above computations together, we can ensure that $M_{n}$ is a supermartingale if the condition $E\leq 1$ is satisfied, where
$\begin{split}E&=\sum_{k=1}^{\infty}45\lambda^{-72\rho}(2\lambda)^{k}(3\lambda)^{-k}\alpha^{-1}+45q\lambda^{-72\rho}+\sum_{k=1}^{\infty}45\lambda^{-72\rho}(4\lambda)^{-k}\left(9\lambda^{2}q\mathbbm{t}\right)(3\lambda)^{k}\alpha\\ &=45\lambda^{-72\rho}\left(\alpha^{-1}\frac{2\lambda}{3\lambda-2\lambda}+q+9\lambda^{2}q\mathbbm{t}\alpha\frac{3\lambda}{4\lambda-3\lambda}\right)\\ &\leq 45\lambda^{-72\rho}\left(\frac{2}{\alpha}+q+27\rho q\alpha\log\frac{1}{\lambda}\right)\end{split}$
We can thus ensure $E\leq 1$ by setting $\alpha=91\lambda^{-72\rho}$, in which case the first term is less than $\frac{90}{91}$ and the other two are $o(1)$ as $\lambda\downarrow 0$. The supermartingale $M_{n}$ then converges a.s. to $M_{\infty}\geq\phi^{\mathcal{V},[v]}_{\infty}(v)$, and as usual by optional stopping we have
$m^{\mathcal{V},[v]}_{\infty}\leq\mathbb{E}\left(M_{\infty}\right)\leq M_{0}=(3\lambda)^{\operatorname{d}(o,v)}\left(91\lambda^{-72\rho}\right)^{n_{\rm big}(o,v)}$
@@
We now have all prerequesite results to upper bound the survival probability on the random tree $\mathcal{A}$.
@@
For any $\delta>0$, the point infection at the root of the power law (truncated at $\mathbbm{t}$) Galton-Watson tree $(\mathcal{A},o)$ will survive with probability less than
@@
$p:=\mathbb{P}\left(D\geq\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor\right)\sim A\left(8\lambda^{2}\right)^{\tau-2},$
which is much smaller than $q$ when $\rho$ is small. It will then be extremely unlikely that it has more than $q\mathbbm{t}$ high-degree neighbours (or $q\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor$ if $d\leq\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor$). In order to bound the survival probability on $\mathcal{A}$ we thus:
  • show that a ball $B(o,k)$ in $\mathcal{A}$ is likely to be a subtree of $\mathcal{B}$.
  • on this event, bound the probability of hitting $\partial B(o,k)$, and thus the survival probability on $\mathcal{A}$, thanks to Lemma 5.4.
More precisely, take
$k:=\left\lceil\lambda^{150\rho-2}\right\rceil\text{ and }\partial B(o,k):=\left\{v\in\mathcal{A}:\operatorname{d}(o,v)=k\right\},$
and let $\mathcal{E}$ be the event that $B(o,k)$ is a subtree of $\mathcal{B}$. Recall in this case that we assume $\mathbbm{t}(\lambda)\leq\rho\frac{\log\frac{1}{\lambda}}{\lambda^{2}}$ and so in the strict interior of this ball we cannot have more than
@@
Now, on the event $\mathcal{E}$, we introduce $\xi^{[\partial B(o,k)]}$ the contact process with absorption at all vertices $v\in\partial B(o,k)$. We can then use Lemma 5.4 to obtain that the mean number of infections hitting any $v\in\partial B(o,k)$ has
$\mathbb{E}\left[\xi^{[\partial B(o,k)]}_{\infty}(v)\right]\leq\left(91\lambda\right)^{\operatorname{d}(o,v)-72\rho\,n_{\rm big}(o,v)}\leq\left(91\lambda\right)^{(1-72\rho)k}.$
Writing $I(\partial B(0,k))=\sum_{v\in\partial B(o,k)}\xi^{[\partial B(o,k)]}_{\infty}(v)$ for the total number of infections hitting $\partial B(0,k)$, we conclude by linearity of the expectation
$\mathbb{E}\left(I(\partial B(o,k))\mathbbm{1}_{\mathcal{E}}\right)\leq\left(91\lambda\right)^{(1-72\rho)k}\mathbb{E}\left(\left|\partial B(o,k)\right|\mathbbm{1}_{\mathcal{E}}\right)\leq\left(91\lambda\right)^{(1-72\rho)k}\mathbb{E}\left(\left|\partial B(o,k)\right|\right).$
It thus remains to control the expected number of targets. The offspring mean when $\tau\in(2.5,3)$ has
$\begin{split}\mathbb{E}(D)&\leq\int_{0}^{\mathbbm{t}-1}\mathbb{P}(D>x){\rm d}x\sim\int_{0}^{\mathbbm{t}}\left(Ax^{2-\tau}-A\mathbbm{t}^{2-\tau}\right){\rm d}x\\ &=A\left(\frac{1}{3-\tau}-1\right)\mathbbm{t}^{3-\tau}\sim A\rho^{3-\tau}\left(\frac{\tau-2}{3-\tau}\right)\lambda^{2\tau-6}\log^{3-\tau}\lambda.\end{split}$
Thus
$\begin{split}\mathbb{E}\left(I(\partial B(o,k))\mathbbm{1}_{\mathcal{E}}\right)&\leq\left(91\lambda\right)^{(1-72\rho)k}\mathbb{E}\left(D\right)^{k}\\ &\leq\exp\left(-\left(2\tau-5-73\rho\right)\lambda^{150\rho-2}\log\frac{1}{\lambda}\right)\end{split}$
where we require small $\rho$ to make a useful bound. Altogether by the union bound
(infection $\begin{split}\mathbb{P}(\text{infection survives})&\leq\mathbb{P}(\mathcal{E}^{c})+\mathbb{P}\left(\{I(\partial B(o,k))>0\}\cap\mathcal{E}\right)\\ &\leq\exp\left(-\frac{1}{10}\lambda^{145\rho-2}\right)+\exp\left(-\left(2\tau-5-73\rho\right)\lambda^{150\rho-2}\log\frac{1}{\lambda}\right)\\ &\leq e^{-\lambda^{-2+\delta}}\end{split}$
@@
Finally, in the excluded $\tau=3$ case we instead calculate $\mathbb{E}(D)\sim 2A\log\tfrac{1}{\lambda}$. This has slower logarithmic growth as $\lambda\downarrow 0$ than the polynomial growth we saw above for any $\tau\in(2.5,3)$, and so by comparison to one of the other $\tau$ values the same bounds still apply. ∎
@@ 6.1 Large truncation threshold
For $\tau>3$, an significant difference with the case $\tau\leq 3$ is that big vertices are (typically) not direct neighbours but further apart. More precisely, we have to look at the neighbourhood of a big vertex up to distance $c\log(1/\lambda)$ for some $c>(2\tau-6)/\log\mu$ in order to typically find other big vertices. The larger threshold $\mathbbm{t}=\rho\lambda^{-2}\log^{2}(1/\lambda)$ however allows for the local survival at a big vertex to last so long that we typically will infect those close big vertices.
Calling “big vertex” a vertex with degree larger than $\mathbbm{t}/2$, the large truncation result in the case $\tau>3$ follows from the following proposition:
@@
For large enough $\rho$ and small $\lambda>0$, the survival probability given the root is big, is bounded below by $1/3$.
We start from a big vertex and look at $D_{R}$ the total amount of vertices at distance exactly $R=\left\lfloor c\log(1/\lambda)\right\rfloor.$ We have
$\operatorname{\mathbb{E}}[D_{R}]\geq\frac{\mathbbm{t}}{2}\mathbb{E}[D]^{R-1},$
@@
Note that $\mathbb{V}{\rm ar}D$ is of order $O(\mathbbm{t}^{4-\tau})=o(\mathbbm{t})$ for $\tau\in(3,4)$, and of logarithmic or constant order if $\tau\geq 4$, thus $\mathbb{V}{\rm ar}\left|\partial B\left(o,R\right)\right|\ll\left(\mathbb{E}\left|\partial B\left(o,R\right)\right|\right)^{2}$ and so by the second moment method we have
@@
Now we consider the infection. By [mountford2013metastable, Lemma 3.1(i)] we have probability $\frac{1}{e}(1-o(1))$ to infect a proportion larger than $\nicefrac{{\lambda}}{{16e}}$ of the neighbours of the root, and from this point we can use [mountford2013metastable, Lemma 3.2] to infect the close big vertices, provided the truncation threshold was large enough. We state here a variation of this lemma directly applicable to our settings:
Lemma 6.2 ( [mountford2013metastable] Lemma 3.2 ).
@@
$\mathbb{P}\left(\exists t>0:\frac{\xi_{t}(B(y,1))}{\lambda\deg(y)}>\frac{1}{16e}\Bigg{|}\frac{\xi_{0}(B(x,1))}{\lambda\deg(x)}>\frac{1}{16e}\right)>1-2e^{-c\log^{2}\frac{1}{\lambda}}.$
Taking $\rho$ large as in this lemma, by the union bound, we then succeed in infecting all the close to big vertices with failure probability bounded by
$O^{\log\frac{1}{\lambda}}\left(e^{-c\log^{2}\frac{1}{\lambda}}\lambda^{(2\tau-6)-c\log\operatorname{\mathbb{E}}[D]}\right)\to 0.$
Finally, from each of these infected big vertices, we can imagine cutting the parent edge and start an independent infection process in its own branch, thus obtaining a branching process of big vertices ever hit by the infection, which is clearly supercritical and has survival probability larger than $\frac{1}{e}(1-o(1))>1/3.$

### v2

@@ 5.2 Small truncation threshold
We let $w\sim o$ be the neighbour of the root on the path from $o$ to $v$. While infected, the root keeps infecting $w$ at rate $\lambda$, so we can bound [^3] Note that the number of infections from $o$ to $w$ has a direct influence on the possible reinfections of $o$ from $w$ and thus on the total time the root is infected. So we do NOT condition on the total time the center is infected to obtain our upper bound on the mean number of infections from $o$ to $w$. the mean total number of infections from $o$ to $w$ by the contact process by $\lambda\cdot 53e^{72\lambda^{2}\mathbbm{d}}\leq 27e^{72\lambda^{2}\mathbbm{d}}(2\lambda)$. This allows to conclude the case $k=1$ where $w=v$. Suppose now $k\geq 2$, and let $(U_{n})_{n\geq 1}$ be the ordered infection times from $o$ to $w$, with $U_{n}=+\infty$ if the number of such infections is less than $n$. As previously explained, we then have $N:=\inf\{n\geq 0,U_{n+1}=+\infty\}$ random with $\operatorname{\mathbb{E}}[N]\leq 27e^{72\lambda^{2}\mathbbm{d}}(2\lambda)$. Now, any infection path from $o$ to $v$ can be decomposed into a first part of the path leading to $w$ at some time $U_{n}$, and a second part of the path from $w$ to $v$ without returning to the root, namely living in the subtree $\mathscr{T}(w)$ obtained by disconnecting $w$ from the root. So the number of infections sent to $v$ is bounded by
$\sum_{n\geq 1}\mathbbm{1}_{U_{n}<+\infty}X_{n}(w,v),$
where $X_{n}(w,v)$ is the number of infections sent to $v$ for the contact process in $\mathscr{T}(w)$ started at time $U_{n}$ with only $w$ infected. Note that $X_{n}(w,v)$ is not independent from $N$, from $U_{n+1}$ or from $X_{n+1}(w,v)$, but still the $U_{n}$ are stopping times, and conditionally on $U_{n}$ being finite, we can bound the expectation of $X_{n}(w,v)$ by $(2\lambda)^{d(w,v)},$ as the tree $\mathscr{T}(w)$ has degrees bounded by $1/8\lambda^{2}$. So the mean number of infections sent from $o$ to $v$ has
$m(o,v)$ $\leq\operatorname{\mathbb{E}}\left[\sum_{n\geq 1}\mathbbm{1}_{U_{n}<+\infty}X_{n}(w,v)\right]$
$\leq\operatorname{\mathbb{E}}[N](2\lambda)^{k-1}\leq 27e^{72\lambda^{2}\mathbbm{d}}(2\lambda)^{k}.$
The former results holds on trees which contains vertices of degree $\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor$ and only one high-degree vertex. On more complicated trees with several high-degree vertices, we will bound the infection by a CP-BRW with sinks all high-degree vertices. However, in order for this CP-BRW to die out locally (and thus for the upper bound to be of any use), we will request a control on the high degrees (have them bounded by $\mathbbm{t}$), as well as on the number of such high degree vertices. More precisely, we consider a tree where only a small proportion $q$ of the neighbours of any given vertex has degree $\mathbbm{t}$.
@@
$m(o,v)\leq\left(55\lambda\right)^{\operatorname{d}(o,v)-72\rho\,n_{\rm big}(o,v)}.$
@@
$m^{\mathcal{V}}(w,u)\leq 27\lambda^{-72\rho}(2\lambda)^{d(w,u)}.$
@@
$27\lambda^{-72\rho}\left((2\lambda)^{k}\cdot q\mathbbm{t}\left(\frac{1}{8\lambda^{2}}\right)^{k-1}+\sum_{m=1}^{\infty}(2\lambda)^{2m+k}\cdot q\left(\frac{1}{8\lambda^{2}}\right)^{m+k}\right)$
$=27\lambda^{-72\rho}(4\lambda)^{-k}\left(q\mathbbm{t}\cdot 8\lambda^{2}+q\right)\leq 45\lambda^{-72\rho}(4\lambda)^{-k}\left(9\lambda^{2}q\mathbbm{t}\right).$
@@
$\sum_{m=1}^{\infty}q\left(\frac{1}{8\lambda^{2}}\right)^{m}\cdot 27\lambda^{-72\rho}(2\lambda)^{2m}=27q\lambda^{-72\rho}.$
@@
Recall now the notation $\left(\phi^{\mathcal{V},[v]}_{n}\right)_{n\in\mathbb{N}}$ from Section 2.4 for the BRW on the set of sinks with absorption at $v$, and define
@@
for some parameter $\alpha$. We can ensure that $M_{n}$ is a supermartingale if the condition $\operatorname{\mathbb{E}}\left[\nicefrac{{M_{n+1}}}{{M_{n}}}\right]\leq 1$ is satisfied, and putting the above computations together gives
$\begin{split}\operatorname{\mathbb{E}}\left[\frac{M_{n+1}}{M_{n}}\right]\leq E&=\sum_{k=1}^{\infty}27\lambda^{-72\rho}(2\lambda)^{k}(3\lambda)^{-k}\alpha^{-1}+27q\lambda^{-72\rho}+\sum_{k=1}^{\infty}27\lambda^{-72\rho}(4\lambda)^{-k}\left(9\lambda^{2}q\mathbbm{t}\right)(3\lambda)^{k}\alpha\\ &=27\lambda^{-72\rho}\left(\alpha^{-1}\frac{2\lambda}{3\lambda-2\lambda}+q+9\lambda^{2}q\mathbbm{t}\alpha\frac{3\lambda}{4\lambda-3\lambda}\right)\\ &\leq 27\lambda^{-72\rho}\left(\frac{2}{\alpha}+q+27\rho q\alpha\log\frac{1}{\lambda}\right)\end{split}$
We can thus ensure $E\leq 1$ by setting $\alpha=55\lambda^{-72\rho}$, in which case the first term is less than $\nicefrac{{54}}{{55}}$ and the other two are $o(1)$ as $\lambda\downarrow 0$. The supermartingale $M_{n}$ then converges a.s. to $M_{\infty}\geq\phi^{\mathcal{V},[v]}_{\infty}(v)$, and as usual by optional stopping we have
$m^{\mathcal{V}}_{\infty}(o,v)\leq\operatorname{\mathbb{E}}_{\mathbbm{1}_{o}}\left[M_{\infty}\right]\leq M_{0}=(3\lambda)^{\operatorname{d}(o,v)}\left(55\lambda^{-72\rho}\right)^{n_{\rm big}(o,v)}$
@@
Write $(\mathcal{A},o)$ for the rooted Galton-Watson tree with power law offspring distribution but truncated at $\mathbbm{t}$. We now have all prerequesite results to upper bound the survival probability on the random tree $\mathcal{A}$.
@@
For any $\delta>0$, the contact process with
$\xi_{0}=\mathbbm{1}_{o}$
on $(\mathcal{A},o)$ will survive to infinity with probability less than
@@
$p:=\mathbb{P}\left(D\geq\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor\right)\sim c_{X}\left(8\lambda^{2}\right)^{\tau-2},$
which is much smaller than $q$ when $\rho$ is small. It will then be extremely unlikely that it has more than $q\mathbbm{t}$ high-degree neighbours (or $q\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor$ if $d\leq\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor$). We now choose a distance $k$ as
$k:=\left\lceil\lambda^{150\rho-2}\right\rceil.$
We consider the ball $B(o,k)$ in $\mathcal{A}$, namely the subtree of $\mathcal{A}$ consisting of all vertices at distance at most $k$ from $o$, as well as its “boundary” $\partial B(o,k)$, which consists of all vertices as distance exactly $k$ from $o$. In order to bound the survival probability on $\mathcal{A}$ we:
  • show that $B(o,k)$ is likely to be a subtree of $\mathcal{B}$. We write $\mathcal{E}$ for this event.
  • on the event $\mathcal{E}$, bound the probability of hitting $\partial B(o,k)$, and thus the survival probability on $\mathcal{A}$, thanks to Lemma 5.4.
Recall in this case that we assume $\mathbbm{t}(\lambda)\leq\rho\frac{\log\frac{1}{\lambda}}{\lambda^{2}}$ and so in the strict interior of this ball we cannot have more than
@@
With the event $\mathcal{E}$ we write, as in the notation of (4),
$m_{\mathcal{E}}(o,\partial B(o,k)):=\sum_{v\in\partial B(o,k)}m_{\mathcal{E}}^{\partial B(o,k)}(o,v)$
for the mean infections sent, restricted to the event $\mathcal{E}$. We can then use Lemma 5.4 to obtain that the mean number of infections hitting any $v\in\partial B(o,k)$ has
$m_{\mathcal{E}}^{\partial B(o,k)}(o,v)\leq\left(55\lambda\right)^{\operatorname{d}(o,v)-72\rho\,n_{\rm big}(o,v)}\leq\left(55\lambda\right)^{(1-72\rho)k}.$
We conclude by linearity of the expectation
$m_{\mathcal{E}}(o,\partial B(o,k))\leq\left(55\lambda\right)^{(1-72\rho)k}\operatorname{\mathbb{E}}\left[\left|\partial B(o,k)\right|\mathbbm{1}_{\mathcal{E}}\right]\leq\left(55\lambda\right)^{(1-72\rho)k}\operatorname{\mathbb{E}}\left[\left|\partial B(o,k)\right|\right].$
It remains to control the expected number of targets. The $\tau\in(2.5,3)$ offspring mean is
$\begin{split}\operatorname{\mathbb{E}}[D]&\leq\int_{0}^{\mathbbm{t}-1}\mathbb{P}(D>x){\rm d}x\sim\int_{0}^{\mathbbm{t}}\left(c_{X}x^{2-\tau}-c_{X}\mathbbm{t}^{2-\tau}\right){\rm d}x\\ &=c_{X}\left(\frac{1}{3-\tau}-1\right)\mathbbm{t}^{3-\tau}\sim c_{X}\rho^{3-\tau}\left(\frac{\tau-2}{3-\tau}\right)\lambda^{2\tau-6}\log^{3-\tau}\lambda.\end{split}$
and so
$\begin{split}m_{\mathcal{E}}(o,\partial B(o,k))&\leq\left(55\lambda\right)^{(1-72\rho)k}\operatorname{\mathbb{E}}\left[D\right]^{k}\\ &\leq\exp\left(-\left(2\tau-5-73\rho\right)\lambda^{150\rho-2}\log\frac{1}{\lambda}\right)\end{split}$
where we require small $\rho$ to make a useful bound. Altogether by the union bound and Markov’s inequality
(infection $\begin{split}\mathbb{P}(\text{infection survives})&\leq\mathbb{P}(\mathcal{E}^{c})+m_{\mathcal{E}}(o,\partial B(o,k))\\ &\leq\exp\left(-\frac{1}{10}\lambda^{145\rho-2}\right)+\exp\left(-\left(2\tau-5-73\rho\right)\lambda^{150\rho-2}\log\frac{1}{\lambda}\right)\\ &\leq e^{-\lambda^{-2+\delta}}\end{split}$
@@
Finally, in the excluded $\tau=3$ case we instead calculate $\operatorname{\mathbb{E}}[D]\sim 2c_{X}\log\tfrac{1}{\lambda}$. This has slower logarithmic growth as $\lambda\downarrow 0$ than the polynomial growth we saw above for any $\tau\in(2.5,3)$, and so by comparison to one of the other $\tau$ values the same bounds still apply. ∎
@@ 6 Power-law exponent $\tau>3$
In this section we address the final parameter regime $\tau>3$ of our main theorem. First in Section 6.1 we lower bound the survival probability in the supercritical case. The upper bound for the subcritical case is addressed in Section 6.2, in part by Lemma 6.3 for the event of the environment and then at the end of the section we argue for the subevent of the process. For this, we introduce the environment offspring mean parameter
$\mu:=\operatorname{\mathbb{E}}(X)>1$
which was infinite in the previous sections.
@@ 6.1 Large truncation threshold
For $\tau>3$, an significant difference with the case $\tau\leq 3$ is that big vertices are (typically) not direct neighbours but further apart. More precisely, we have to look at the neighbourhood of a big vertex up to distance $c\log(\nicefrac{{1}}{{\lambda}})$ for some $c>(2\tau-6)/\log\mu$ in order to typically find other big vertices. The larger threshold $\mathbbm{t}=\rho\lambda^{-2}\log^{2}(\nicefrac{{1}}{{\lambda}})$ however allows for the local survival at a big vertex to last so long that we typically will infect those close big vertices.
Calling “big vertex” a vertex with degree larger than $\nicefrac{{\mathbbm{t}}}{{2}}$, the large truncation result in the case $\tau>3$ follows from the following proposition:
@@
For large enough $\rho$ and small $\lambda>0$, the survival probability given the root is big, is bounded below by $\nicefrac{{1}}{{3}}$.
We start from a big vertex and look at $D_{R}$ the total amount of vertices at distance exactly $R=\left\lfloor c\log(\nicefrac{{1}}{{\lambda}})\right\rfloor.$ We have
$\operatorname{\mathbb{E}}[D_{R}]\geq\frac{\mathbbm{t}}{2}\operatorname{\mathbb{E}}[D]^{R-1},$
@@
Note that $\mathbb{V}{\rm ar}D$ is of order $O(\mathbbm{t}^{4-\tau})=o(\mathbbm{t})$ for $\tau\in(3,4)$, and of logarithmic or constant order if $\tau\geq 4$, thus $\mathbb{V}{\rm ar}\left|\partial B\left(o,R\right)\right|\ll\left(\operatorname{\mathbb{E}}\left|\partial B\left(o,R\right)\right|\right)^{2}$ and so by the second moment method we have
@@
Now we consider the infection. By [mountford2013metastable, Lemma 3.1(i)] we have probability $\nicefrac{{1}}{{e}}-o(1)$ to infect a proportion larger than ${\lambda}/{16e}$ of the neighbours of the root, and from this point we can use the following lemma to infect all nearby vertices, provided the truncation threshold was large enough. We state here a variation of this lemma directly applicable to our settings:
Lemma 6.2 ( [mountford2013metastable, Lemma 3.2] ).
@@
$\mathbb{P}\left(\exists t>0:\frac{\xi_{t}(B(y,1))}{\lambda\deg(y)}>\frac{1}{16e}\Bigg{|}\frac{\xi_{0}(B(x,1))}{\lambda\deg(x)}>\frac{1}{16e}\right)>1-2e^{-c\log^{2}\frac{1}{\lambda}}$
where $\xi_{t}$ applied to a set of vertices denotes the total number of infected vertices in that set.
Taking $\rho$ large as in this Lemma, by the union bound, we then succeed in infecting all the nearby vertices $y$ of the Lemma with failure probability bounded by
$1-\left(1-2e^{-c\log^{2}\frac{1}{\lambda}}\right)^{|B(x,c\log\nicefrac{{1}}{{\lambda}})|}=O^{\log\frac{1}{\lambda}}\left(e^{-c\log^{2}\frac{1}{\lambda}}\lambda^{(2\tau-6)-c\log\operatorname{\mathbb{E}}[D]}\right)\to 0$
Altogether, we thus have probability $\nicefrac{{1}}{{e}}-o(1)$ to infect all nearby vertices, and in particular all the big vertices at distance $R$ from the root. Now, from each of these infected big vertices, we can imagine cutting the parent edge and start an independent infection process in its own branch, thus obtaining a branching process of big vertices ever hit by the infection, which has an offspring mean tending to infinity and so is clearly supercritical, with survival probability larger than $\nicefrac{{1}}{{e}}-o(1)>\nicefrac{{1}}{{3}}.$